OpenStudy (ammarah):

Use the 45-45-90 Triangle Theorem to find the sine and cosine of a 45 angle.....

OpenStudy (aravindg):

well the 45-45-90 triangle will have 2 sides equal and the other will be the hypotenuse

OpenStudy (anonymous):

a 45-45-90 triangle is isoceles triangle

OpenStudy (aravindg):

so take the equal sides as 1 then hypotenuse will be root(1^2+1^2)=root 2

OpenStudy (aravindg):

cos 45 =adj side/hypotenuse =1/root 2

OpenStudy (aravindg):

\[hence\ \cos 45=\frac{1}{\sqrt{2}}\]